# Fast and Accurate Task Planning using Neuro-Symbolic Language Models and Multi-level Goal Decomposition

### v1

@@ V-B Performance Analysis
Comparisons: The CoT planner is the fastest among the four planners but has the lowest accuracy, with the success rate approaching nearly zero as $n$ increases; on the other hand, the success rate for the FD planner is always 100%, but its planning time increases exponentially as $n$ grows. This indicates that both baseline methods struggle to solve long-sequential problems in highly complex search spaces. In contrast, our symbolic LLM planner consistently achieved a success rate of 100%, and our MCTS planner obtained on average $98.5\%,92.6\%,88.2\%$ success rates for Barman-new, Blocksworld-new and Gripper-new domains, respectively. Compared to the CoT Planner, on average, the planning times of our symbolic and MCTS planners are 6.5×6.5$\times$6.5 ×/3.8×3.8$\times$3.8 × (Barman-new), 4.9×4.9$\times$4.9 ×/10.2×10.2$\times$10.2 × (Blocksworld-new), and 3.36×3.36$\times$3.36 ×/8×8$\times$8 × (Gripper-new) slower, respectively.
@@
It is generally difficult to compare the performance of our method against other state-of-the-art, LLM-based methods since they use different LLMs or generate non-deterministic results. However, one can estimate comparisons based on the original authors’ report. [13] show very low success rates (almost zero) for complex benchmarks like ours, [17] show similar success rates like ours for the single gripper benchmark whereas ours is dual, more complex setup, and [18] show slightly inferior success rates than ours for the Blocksworld when $n\leq 4$, but it is unclear how it would perform for $n>4$. Thus, even though this comparative study is not purely experimental, one can say that the performance of our methods is substantially better than the state of the art.
@@
Symbolc LLM vs. MCTS LLM: In the Barman-new domain, where the MDL between each subgoal (making each cocktail) is long, and the domain’s state space $\mathcal{S}$ is extensive, the planning time for the symbolic LLM Planner increases rapidly as $n$ grows. In contrast, the MCTS LLM planner exhibits an almost linear increase in planning time in terms of $n$, resulting in faster performance than the symbolic LLM planner. However, in the Blocksworld-new and Gripper-new domains, the planning time for the symbolic LLM planner does not increase as quickly as in the Barman-new domain, and it was faster than the MCTS LLM planner, probably because those domains are less complex than the Barman-new domain, and the MDL between subgoals is moderate.
@@
Sampled Plans: We performed further experiments on the number of sampled plans used by the MCTS LLM planner by varying $3\leq n_{s}\leq 5$, and observed a general trend of higher success rates, accompanied by an increase in planning time. As noted in [5], there seems to be a limit to success rate improvement when increasing $n_{s}$, as the complexity of the search space has an upper bound.
@@
Ablation Study on Goal Decomposition: We conducted an ablation study on the effectiveness of goal decomposition. We execute our MCTS LLM planner with $n_{s}=5$ with and without goal decompositions. As shown in Fig. 3, our planner with goal decomposition achieved a much higher success rate than the one without it, whereas the planner without goal decomposition approached zero success rates for complex problems.
@@ V-C Robot Demonstration
We successfully conducted planning experiments with a real robot in the Blocksworld-new domain to demonstrate the practicality of our neuro-symbolic robot task planners. For the real robot demonstration, we used dual UR5e manipulators equipped with Robotiq 3F grippers. An Intel RealSense D455 RGBD camera was employed for visual input, fixed above the table to provide a top-down view. Also, for the Barman-new domain, we conducted experiments in the CoppeliaSim environment [30], which was set up similarly to the real robot setup described above. For both experiments, our task planners are integrated into the MoveIt motion planner[31] in ROS via the translated action primitives. Key robot action primitives, such as pick and place, were predefined using MoveIt, and the task planning results were converted into code composed of these action primitives using the LLM. Once executed, the corresponding robot actions are carried out accordingly.
@@ VI CONCLUSION AND FUTURE WORK
This paper proposes a novel task-planning method based on neuro-symbolic language models by decomposing a complicated, long-sequential goal into multi-level subgoals. Our planner performs much faster than the baseline symbolic methods, achieving high accuracy. We would like to pursue a couple of future directions to improve our current method. The criterion for choosing the level of goal decompositions and picking either a symbolic or MCTS planner for the subgoal is empirical. One needs more systematic or thorough studies of the decision problem. We also need more investigation into integrating our tasking planning pipeline to motion planning, i.e., the TAMP.

### v2

@@ V-B Performance Analysis
Comparisons: The CoT planner is the fastest among the four but has the lowest accuracy, with its success rate approaching nearly zero as $n$ increases; on the other hand, the FD planner maintains a 100% success rate, but its planning time increases exponentially as $n$ grows. This indicates that both baseline methods struggle to solve long-sequential problems in highly complex search spaces. In contrast, our Symbolic LLM planner consistently achieved a success rate of 100%, and our MCTS planner obtained on average $98.5\%,92.6\%,88.2\%$ success rates for Barman-new, Blocksworld-new, and Gripper-new domains, respectively. Compared to the CoT Planner, on average, the planning times of our symbolic and MCTS planners are 6.5×6.5$\times$6.5 ×/3.8×3.8$\times$3.8 × (Barman-new), 4.9×4.9$\times$4.9 ×/10.2×10.2$\times$10.2 × (Blocksworld-new), and 3.36×3.36$\times$3.36 ×/8×8$\times$8 × (Gripper-new) slower.
@@
It is difficult to compare the performance of our method against other state-of-the-art, LLM-based methods since they use different LLMs or generate non-deterministic results. However, one can estimate comparisons based on the original authors’ report. [15] show very low success rates (almost zero) for complex benchmarks like ours, [19] show similar success rates like ours for the single robot benchmark whereas ours is multiple, more complex setup, and [20] show slightly inferior success rates than ours for the Blocksworld when $n\leq 4$, but it is unclear how it would perform for $n>4$. Even though this comparative study is not purely experimental, one can say that the performance of our methods is substantially better than the existing methods.
@@
Symbolic LLM vs. MCTS LLM: In the Barman-new domain, where each sub-task (making a cocktail) requires a long MDL and the domain’s state space $\mathcal{S}$ is large, the planning time for the Symbolic LLM planner increases rapidly as $n$ grows. In contrast, the MCTS LLM planner shows an almost linear growth in planning time with respect to $n$, resulting in faster performance than the Symbolic LLM planner. However, in the Blocksworld-new and Gripper-new domains, the planning time for the Symbolic LLM planner does not increase as quickly as in the Barman-new domain, and it was faster than the MCTS LLM planner. This is probably because these domains are less complex than the Barman-new domain, with a shorter MDL between subgoals.
@@
Sampled Plans: We performed further experiments on the number of sampled plans used by the MCTS LLM planner by varying $3\leq n_{s}\leq 5$, and observed a general trend of higher success rates, accompanied by an increase in planning time. However, as noted in [5], success rate improvement is limited when $n_{s}$ exceeds a certain point due to the upper bound on search space complexity, with subgoal decomposition further restricting the space in our case.
@@
Ablation Study on Goal Decomposition: We conducted an ablation study on the effectiveness of goal decomposition. We executed our MCTS LLM planner with $n_{s}=5$ with and without goal decompositions. As shown in Fig. 3, our planner with goal decomposition achieved a much higher success rate than the one without it, whereas the planner without goal decomposition approached zero success rates for complex problems.
@@ V-C Robot Demonstration
We conducted planning experiments with a real robot in the Blocksworld-new domain to demonstrate the practicality of our neuro-symbolic robot task planners. For the real robot demonstration, we used dual UR5e manipulators with Robotiq 3F grippers. An Intel RealSense D455 RGBD camera was employed for visual input, fixed above the table for a top-down view. For the Barman-new domain, we conducted experiments in the CoppeliaSim environment [32], which was set up similarly to the real robot setup. For both experiments, our task planners were integrated into the MoveIt motion planner[33] in ROS via the translated action primitives. Key robot action primitives, such as pick and place, were predefined using MoveIt, and task planning results were converted into code composed of these action primitives using the LLM. Once executed, corresponding robot actions were carried out accordingly.
However, we assume that every high-level action primitive has a feasible low-level solution without explicitly handling motion planning or execution failures. This simplification may lead to uncertainties in real-world execution, affecting execution robustness. Addressing these issues in future work could improve overall system reliability.
V-D Failure Analysis
In both real-world and simulation experiments, failures fell into two categories: execution and planning. Execution failures mostly stemmed from stability issues, such as occlusion in cluttered environments leading to inaccurate planning formulation, failed grasps, or collapsed stacks of blocks. Planning failures were more common in the MCTS LLM planner. In the Blocksworld-new domain, it struggled with spatial reasoning and misordered block stacking sequences, while in the Gripper-new domain, it misinterpreted the goal state, occasionally moving irrelevant balls.
@@ VI CONCLUSION AND FUTURE WORK
This paper proposes a novel task-planning method based on neuro-symbolic language models by decomposing a complicated, long-sequential goal into multi-level subgoals. Our planner performs much faster than the baseline symbolic methods, achieving high accuracy. Future improvements include developing automated strategies for selecting the level of goal decomposition and choosing between symbolic and MCTS planners, which currently rely on empirical criteria. Further integration of our task-planning pipeline with motion planning (i.e., the TAMP) is also needed. Moreover, ensuring the generalizability of the subgoal decomposition in more complex real-world tasks remains essential. Expanding evaluations to diverse environments, such as multi-agent systems, will further enhance their robustness and generalizability
ACKNOWLEDGMENT
This work was supported in part by the ITRC/IITP Program (IITP-2025-RS-2020-II201460), and in part by the NRF (NRF-2022R1A2B5B03001385) in South Korea.